# Quantum Natural Gradient optimizer on noisy platforms: QAOA as a case study

### v1

@@ I Introduction
This study will focus on a paradigmatic problem: finding the ground state preparation of the Transverse Field Ising Model (TFIM). In particular, we aim to combine the Quantum Approximate Optimization Algorithm (QAOA) with the Quantum Natural Gradient (QNG) optimizer and evaluate its performance on simulations of realistic noisy quantum devices. The Quantum Natural Gradient was originally proposed in [16] to enhance the optimization of variational quantum algorithms. Several studies have since demonstrated the potential advantages of using QNG techniques in these algorithms [17, 18].
@@ II.1 QAOA Circuit
    Here, $H_{c}$ is the cost Hamiltonian and $H_{m}$ is the mixer Hamiltonian. To explore the whole Hilbert space, it is necessary to have $[H_{c},H_{m}]\neq 0$. Then, the mixer Hamiltonian is chosen as $H_{m}=\sum_{i=1}^{N}\sigma^{x}_{i}$, where $\sigma^{x}_{i}$ are Pauli-X operators acting on the $i$-th qubit.
@@ II.2 Application to the Transverse Field Ising Model
$H_{c}=-J\sum_{i=1}^{N}\sigma^{z}_{i}\sigma^{z}_{i+1}-h\sum_{i=1}^{N}\sigma^{x}_{i}$ (2)
@@
As previously mentioned, the QAOA algorithm follows a sequence of well-defined steps, though it permits flexibility in the choice of Hamiltonians used for evolving the system. Specifically, for the TFIM, it can be demonstrated using exact diagonalization techniques provided by the QuTiP library [21] that choosing the cost Hamiltonian unitary as $U_{zz}(\gamma)=e^{-i\gamma H_{zz}}$ (while keeping the mixer Hamiltonian unitary as the standard form, i.e., $U_{m}(\beta)=e^{-i\beta H_{m}}$), where $H_{zz}=\sum_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}$, guarantees that the ground state is reached at a circuit depth of $P=\left\lfloor\frac{N}{2}\right\rfloor.$
This implies that circuits with such a depth are sufficiently expressive to capture the ground state while maintaining the symmetries of the Hamiltonian. Hence, we adopt this parameterization for the algorithm’s evolution, subject to validation during implementation on the selected platform, which inevitably includes certain approximations, both with and without noise.
The energy value obtained at the end of the optimization procedure $E_{c}$ can be compered with the exact one $E_{0}$, allowing us to compute the accuracy:
$\delta E:=\frac{E_{c}-E_{0}}{\absolutevalue{E_{0}}}.$ (3)
@@ II.3 Quantum Natural Gradient
where $|\psi(\boldsymbol{\theta})\rangle$ is a parameterized quantum state, and we use the shorthand $\partial_{i}=\partial/\partial\theta_{i}$.
@@
$F_{ab}=$ $4\Bigg{[}\langle\psi_{a-1}|H_{a}\Bigg{(}\prod_{k=a}^{b-1}e^{i\theta_{k}H_{k}}\Bigg{)}H_{b}|\psi_{b-1}\rangle-$
@@
This expression stems from the unitary parameter embedding defined by QAOA. In this framework, the Hamiltonian $H_{a}=H_{zz}$ when $a$ is odd and $H_{a}=H_{m}$ when $a$ is even, while |$\ket{\psi_{a-1}}$ represents the state obtained after evolving the initial state with $a-1$ parameters.
@@
$F_{aa}=4\left[\langle\psi_{a-1}|H_{a}^{2}|\psi_{a-1}\rangle-\langle\psi_{a-1}|H_{a}|\psi_{a-1}\rangle^{2}\right].$ (16)
@@
$F_{ab}$ $=4\Big{[}\text{Tr}\left(\rho_{a-1}H_{a}\Bigg{(}\prod_{k=a}^{b-1}e^{i\theta_{k}H_{k}}\Bigg{)}H_{b}\rho_{b-1}\right)$
@@
$F_{aa}=4\left[\text{Tr}(\rho_{a-1}H_{a}^{2})-\left(\text{Tr}(\rho_{a-1}H_{a})\right)^{2}\right].$ (18)
@@ III Rydberg atoms implementation
This work focuses on the simulation of pulse sequences with noise and errors using the pulser library [38, 39]. These factors are intrinsic to real-world quantum devices and must be accurately replicated in simulations to devise effective strategies for addressing them. A comprehensive explanation of these phenomena in Rydberg platforms is provided in [5]. Here, we will briefly outline the aspects considered in this study.
@@ IV.1 Ground state preparation
As explained in the previous sections, the QAOA algorithm adheres to a specific protocol, which in our case involves initializing the state in |$\ket{+}^{\otimes N}$ and evolving it using the Hamiltonians $H_{zz}$ and $H_{m}$. By strictly following this procedure, there is numerical evidence that the ground state can be achieved when $P=\lfloor N/2\rfloor$ [6].
@@
We aimed at exploring this aspect by plotting how fidelity varies with the number of layers $P$ for different system sizes, both in the presence and absence of noise (see Fig. 3). We opted to present the fidelity trends because, in some cases, final states with nearly identical energies may still exhibit low fidelity values. As our primary goal is state preparation, our focus is on how closely the achieved state overlaps with the true ground state. The fidelity between a mixed state and a pure state is defined as $F=\bra{\psi_{g}}\rho\ket{\psi_{g}}$, where $\psi_{g}$ is the ground state vector determined through exact diagonalization. According to the simulation results, it is clear that in the noiseless scenario, the ground state is achieved with an accuracy $\delta E_{opt}<10^{-9}$ when $P=\lfloor N/2\rfloor+1$, where $\delta E_{\text{opt}}$ represents the best accuracy obtained after optimization. This accuracy is calculated as the difference between the energy value obtained by the algorithm at convergence and the true value (see Eq. (3)).
@@ IV.2.1 Noiseless Case
Figure 5: Comparison of noiseless accuracies for the Vanilla, QNG diagonal approximation, and QNG full matrix methods as a function of system size $N$. Each panel displays the distribution of 50 accuracy values (measured as $\delta E$) for each $N$, obtained from different runs, with the median value of accuracies represented by solid lines. The most populated accuracy sector is at $10^{-10}$, indicating successful convergence to the ground state, especially for the QNG Full Matrix which is more consistent as $N$ increases.
@@ IV.2.2 Laser noise and SPAM errors cases
Fig. 7 provides further insights by presenting the accuracy distribution $\delta E$ across 50 runs for the same three methods. For laser noise (top row), the Vanilla method (left) shows a sharp increase in $\delta E$ with $N$, along with a wide distribution spanning several orders of magnitude. This indicates its pronounced sensitivity to noise and poor reliability. In contrast, the QNG diagonal approximation (center) reduces variance and the median is slightly below compered to the Vanilla one, reflecting enhanced robustness to laser noise. The QNG full matrix method (right) delivers superior performance, maintaining low $\delta E$ values with minimal spreadness of them, demonstrating very good stability and accuracy.
@@ V.2 Noise Model Derived from IBM Device Calibration Data
To simulate the realistic behavior of quantum circuits on physical hardware, we incorporated a noise model based on ibm_sherbrooke’s calibration data. Using Qiskit’s Aer simulator [41], we utilized the backend properties of ibm_sherbrooke to define noise characteristics, which were then applied during circuit execution. This noise model accounts for several key factors:
@@ VI.2 Comparison of QNG against Vanilla gradient descent
We compared the performance of the QNG optimizer, using both the full Fubini-Study metric and its diagonal approximation, against Vanilla Gradient Descent in both noiseless and noisy scenarios.
@@ VII Conclusions and outlook
In summary, the QNG optimizer is a highly advantageous method for enhancing the optimization of variational algorithms. It enables faster average convergence, greater robustness against random parameter initializations, and an improved ability to avoid certain local minima. Furthermore, our results demonstrate that QNG exhibits robustness against realistic noise during optimization. Thus, we firmly believe that incorporating QNG for the optimization of variational quantum algorithms on NISQ devices could prove to be a highly valuable tool.

### v2

@@ I Introduction
This study will focus on a paradigmatic problem: finding the ground state preparation of the Transverse Field Ising Model (TFIM). In particular, we aim to combine the Quantum Approximate Optimization Algorithm (QAOA) with the Quantum Natural Gradient (QNG) optimizer and evaluate its performance on simulations of realistic noisy quantum devices. The Quantum Natural Gradient was originally proposed in [16] to enhance the optimization of variational quantum algorithms. Several studies have since demonstrated the potential advantages of using QNG techniques in these algorithms [17, 18, 19].
@@
In the current NISQ era, gradient-free optimization techniques are often preferred for QAOA due to their lower measurement overhead and simpler implementation. In this work, we deliberately focus on gradient-based methods—not as a recommendation for near-term optimization, but rather as a theoretical and exploratory benchmark to assess the robustness of QNG under realistic noise models.
Indeed, the primary objective of this work is not to demonstrate state-of-the-art optimization performance, but rather to investigate how the QNG responds to realistic noise conditions typical of current quantum platforms. For this reason, we adopt the TFIM as a toy model, serving as a well-understood and controllable testbed for benchmarking the behavior of QNG.
@@ II.1 QAOA Circuit
    Here, $H_{c}$ is the cost Hamiltonian and $H_{m}$ is the mixer Hamiltonian. To explore the whole Hilbert space, it is necessary to have $H_{c},H_{m}]\neq 0$. Then, the mixer Hamiltonian is chosen as $H_{m}=\sum_{i=1}^{N}\sigma^{x}_{i}$, where $\sigma^{x}_{i}$ are Pauli-X operators acting on the $i$-th qubit.
@@ II.2 Application to the Transverse Field Ising Model
$H=-J\sum_{i=1}^{N}\sigma^{z}_{i}\sigma^{z}_{i+1}-h\sum_{i=1}^{N}\sigma^{x}_{i}$ (2)
@@
As previously mentioned, the Quantum Approximate Optimization Algorithm (QAOA) follows a sequence of well-defined steps while allowing flexibility in the choice of Hamiltonians used to evolve the system. In particular, for the Transverse Field Ising Model (TFIM), we modify the standard structure of QAOA by defining the unitary operators governing the evolution as $U_{zz}(\gamma)=e^{-i\gamma H_{zz}}$, where $H_{zz}=-J\sum_{i=1}^{N}\sigma^{z}_{i}\sigma^{z}_{i+1}$, for the first evolution of the state. As the mixing Hamiltonian, we use the standard form $U_{m}(\beta)=e^{-i\beta H_{m}}$, where $H_{m}=-h\sum_{i=1}^{N}\sigma^{x}_{i}$. For our numerical simulations, we set $J=1$ and $h=0.5$. However, as the cost function, we measure the expectation value of the full Hamiltonian given in Eq. 2.
This modification is motivated by the fact that the conventional QAOA structure typically employs a mixing Hamiltonian that has the same form as the second term in the problem Hamiltonian (Eq. 2). By adjusting the standard procedure, we eliminate this redundancy.
By implementing the algorithm using the QuTiP library [22], we verified that the ground state is reached at a circuit depth of $P=\left\lfloor\frac{N}{2}\right\rfloor$, by testing different values of $P$ using the quantum natural gradient as optimizer. This preliminary check will also be performed in the implementations based on Rydberg atoms and IBM superconducting qubits, using the Pulser and PennyLane backends, respectively.
This result implies that circuits with such a depth are sufficiently expressive to capture the ground state while preserving the symmetries of the Hamiltonian. Hence, we adopt this parameterization for the algorithm’s evolution, subject to validation during implementation on the selected platform, which inevitably introduces certain approximations, both in the presence and absence of noise.
The energy value obtained at the end of the optimization procedure $E_{c}$ can be compared with the exact one $E_{0}$, allowing us to compute the accuracy:
$|.\delta E:=\frac{E_{c}-E_{0}}{\absolutevalue{E_{0}}}$. (3)
@@ II.3 Quantum Natural Gradient
where $|\psi(\boldsymbol{\theta})\rangle$ is a parameterized quantum state, and we use the shorthand $\partial_{i}\equiv\frac{\partial}{\partial\theta_{i}}$.
@@
$F_{ab}=$ $4\,\text{Re}\Bigg{[}\langle\psi_{a-1}|H_{a}\Bigg{(}\prod_{k=a}^{b-1}e^{i\theta_{k}H_{k}}\Bigg{)}H_{b}|\psi_{b-1}\rangle$-
@@
[…]
@@
$F_{aa}=4\,\text{Re}\left[\langle\psi_{a-1}|H_{a}^{2}|\psi_{a-1}\rangle-\langle\psi_{a-1}|H_{a}|\psi_{a-1}\rangle^{2}\right].$ (16)
@@
$F_{ab}$ $=4\,\text{Re}\Big{[}\text{Tr}\left(\rho_{a-1}H_{a}\Bigg{(}\prod_{k=a}^{b-1}e^{i\theta_{k}H_{k}}\Bigg{)}H_{b}\rho_{b-1}\right)$
@@
$F_{aa}=4\,\text{Re}\left[\text{Tr}(\rho_{a-1}H_{a}^{2})-\left(\text{Tr}(\rho_{a-1}H_{a})\right)^{2}\right].$ (18)
@@
II.4 Computation cost of QNG for QAOA
The cost of computing the diagonal approximation of the Quantum Fisher Information Matrix (QFIM) for QAOA, as defined in Eq. (16), depends on the number of expectation values required for each entry. Each expectation value generally requires a separate quantum circuit preparation. However, this cost can be reduced by grouping observables into sets of commuting operators. Observables within the same commuting group can be measured simultaneously using a single quantum circuit, thereby reducing the total number of distinct circuits required. This technique is a standard practice in the quantum computing community to reduce the resource overhead when estimating observables composed of many terms, such as Hamiltonians [36].
In the worst-case scenario, for a Hamiltonian consisting of $m$ non-commuting terms—where neither the terms of the Hamiltonian nor those of its square commute with each other—one must evaluate $\mathcal{O}(m^{2})$ expectation values to compute a single diagonal entry of the QFIM. This requires preparing and running $\mathcal{O}(m^{2})$ different quantum circuits. To estimate each expectation value with a precision of $\mathcal{O}(\epsilon)$, $\mathcal{O}(1/\epsilon^{2})$ measurements are needed. Consequently, the total measurement cost for estimating one diagonal QFIM entry is $\mathcal{O}(m^{2}/\epsilon^{2})$.
If the Hamiltonians involved in QAOA have different numbers of terms, the cost of estimating the full diagonal QFIM becomes $\mathcal{O}(P(m^{2}+m^{\prime 2})/\epsilon^{2})$, where $m$ and $m^{\prime}$ are the numbers of non-commuting terms in the cost and mixing Hamiltonians, respectively, and $P$ is the number of QAOA layers. However, as previously noted, this cost can be significantly reduced by grouping commuting observables and measuring them using the same circuit. This strategy is also employed in [16] for estimating the block-diagonal approximation of the QFIM.
In our specific case of the transverse field Ising model (TFIM), the cost of computing the diagonal entries of the QFIM is substantially reduced due to the structure of the Hamiltonian. All terms in $H_{zz}$ commute with each other, and the same holds for all terms in $H_{x}$. This allows us to compute all required expectation values for each diagonal QFIM entry using only one quantum state preparation per group. As a result, only $2P$ distinct quantum circuit preparations are needed for the diagonal approximation of the QFIM.
In contrast, computing the gradient requires $4mP$ different quantum circuits, where $m$ is the number of terms in both the cost and mixing Hamiltonians.
On the other hand, computing the full QFIM is substantially more costly because it requires estimating off-diagonal terms between different layers. In general, estimating these terms necessitates the use of the Hadamard test, which requires an auxiliary qubit for each term involved in Eq. (15). To compute a single off-diagonal QFIM entry, one must perform $\mathcal{O}(mm^{\prime})$ different Hadamard tests, each requiring a separate quantum circuit with an auxiliary qubit, where $m$ and $m^{\prime}$ are the number of terms in the respective Hamiltonians $H_{a}$ and $H_{b}$ from Eq. (15). Each such estimate with precision $\mathcal{O}(\epsilon)$ requires $\mathcal{O}(1/\epsilon^{2})$ measurements, leading to a total cost of $\mathcal{O}(mm^{\prime}/\epsilon^{2})$ measurements per off-diagonal entry.
Because the QFIM is symmetric, and assuming $2P$ parameters in the QAOA circuit, there are only $P(2P+1)$ independent entries. Thus, the total cost of estimating the full QFIM in terms of measurements is upper-bounded by $\mathcal{O}(P^{2}m^{2}/\epsilon^{2})$, assuming $m>m^{\prime}$. If $m^{\prime}>m$, the roles can be swapped accordingly.
Importantly, the QFIM must be estimated at each step of the optimization process. Due to this unfavorable scaling, computing the full QFIM becomes impractical for large problem instances. By contrast, the diagonal approximation provides a more feasible alternative, offering potential improvements at a measurement cost similar to that of gradient estimation.
Nonetheless, gradient-based optimization remains out of reach for current near-term quantum devices. There is ongoing research aimed at reducing the complexity of estimating both gradients [37, 38] and the QFIM [39], which may enable these optimization strategies to become viable in practice in the near future.
Additionally, there has been rigorous analysis of the measurement overhead associated with variational quantum algorithms that utilize the QNG for optimization. In [40], the authors derive upper bounds on the number of measurements required to perform QNG at each optimization step for a given target precision. Notably, they show that, asymptotically, the cost of computing the QFIM becomes negligible compared to that of estimating the gradient, particularly as the number of iterations and qubits increases. In this asymptotic regime, QNG introduces only marginal additional overhead relative to standard Gradient Descent, while simultaneously improving convergence speed and overall performance. Furthermore, the authors propose a general strategy for optimally allocating measurement resources between the QFIM and the gradient to minimize total overhead.
@@ III Rydberg atoms implementation
This work focuses on the simulation of pulse sequences (including noise and imperfections) using the Pulser library [44, 45], a Python-based library specifically designed to simulate the dynamics of Rydberg atom systems by numerically solving the full time-dependent Schrödinger equation in the Hilbert space of the system (i.e. full state evolution). These factors are intrinsic to real-world quantum devices and must be accurately replicated in simulations to devise effective strategies for addressing them. A comprehensive explanation of these phenomena in Rydberg platforms is provided in [5]. Here, we will briefly outline the aspects considered in this study.
@@ IV.1 Ground state preparation
[…]
@@
[…]
@@ IV.2.1 Noiseless Case
Figure 5: Comparison of noiseless accuracies for the Vanilla, QNG diagonal approximation, and QNG full matrix methods as a function of system size $N$. Each panel displays the distribution of 50 accuracy values (measured as $\delta E$) for each NN, obtained from different runs, with the median value of accuracies represented by solid lines. The most populated accuracy sector is at $10^{-10}$, indicating successful convergence to the ground state, especially for the QNG Full Matrix which is more consistent as $N$ increases.
@@ IV.2.2 Laser noise and SPAM errors cases
Fig. 7 provides further insights by presenting the accuracy distribution $\delta E$ across 50 runs for the same three methods. For laser noise (top row), the Vanilla method (left) shows a sharp increase in $\delta E$ with $N$, along with a wide distribution spanning several orders of magnitude. This indicates its pronounced sensitivity to noise and poor reliability. In contrast, the QNG diagonal approximation (center) reduces variance and the median is slightly below compared to the Vanilla one, reflecting enhanced robustness to laser noise. The QNG full matrix method (right) delivers superior performance, maintaining low $\delta E$ values with minimal spreadness of them, demonstrating very good stability and accuracy.
@@ V.2 Noise Model Derived from IBM Device Calibration Data
To simulate the realistic behavior of quantum circuits on physical hardware, we incorporated a noise model based on ibm_sherbrooke’s calibration data. Using Qiskit’s Aer simulator [47], we utilized the backend properties of ibm_sherbrooke to define noise characteristics, which were then applied during circuit execution. The calibration data for this specific device is provided in Appendix A, in Table 1. This noise model accounts for several key factors:
@@ VI.2 Comparison of QNG against Vanilla gradient descent
We compared the performance of the QNG optimizer, using both the full Fubini-Study metric and its diagonal approximation, against Vanilla Gradient Descent in both noiseless (see Fig. 10) and noisy scenarios.
@@ VII Conclusions and outlook
In summary, the QNG optimizer is a highly advantageous method for enhancing the optimization of variational algorithms. It enables faster average convergence, greater robustness against random parameter initializations, and an improved ability to avoid certain local minima. Furthermore, our results demonstrate that QNG exhibits robustness against realistic noise during optimization. Thus, we firmly believe that incorporating QNG for the optimization of variational quantum algorithms is highly promising, provided that significant progress is made in reducing the computational cost of gradient evaluation and QFIM estimation on real hardware.
Finally, we want to emphasize that, despite the strong potential of QNG—offering more stable convergence and improved resilience to noise—these benefits come with significant computational costs. In particular, evaluating the QFIM, whether exactly or approximately, introduces substantial overhead in both classical post-processing and quantum measurement resources. As a result, the practical use of QNG is currently limited to small-scale simulations or hypothetical future scenarios where these costs may become more manageable.
VIII Acknowledgments
This work was supported by the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+, the Severo Ochoa Grant CEX2023-001292-S, Generalitat Valenciana grant CIPROM/2022/66, the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). This project has also received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement CaLIGOLA MSCA-2021-SE-01-101086123. RGL is funded by grant CIACIF/2021/136 from Generalitat Valenciana. The work was also partially supported by INFN through the project QUANTUM.
Appendix A Calibration data from the IBM device ibm_sherbrooke
In this Appendix, we present the calibration data from the IBM device ibm_sherbrooke, which were used as parameters for our noise model during circuit execution. The calibration data is shown in Table 1.
Qubit T1 (us) T2 (us) Prob meas0 prep1 Prob meas1 prep0 Readout assignment error ID error RZ error SX error X error ECR error Gate time (ns)
0 351.63 252.89 0.26 0.23333 0.24667 0.0001205 0 0.0001205 0.0001205
1 391.24 383.13 0.0172 0.0228 0.02 0.0001565 0 0.0001565 0.0001565 ecr1_2:0.0049;ecr1_0:0.0064 ecr1_2:533.333;ecr1_0:533.333
2 268.69 170.39 0.0194 0.0338 0.0266 0.0002395 0 0.0002395 0.0002395
3 281.24 259.83 0.53667 0.35 0.44333 0.0002122 0 0.0002122 0.0002122 ecr3_2:0.00619 ecr3_2:533.333
4 410.16 440.13 0.65667 0.2 0.42833 0.0001117 0 0.0001117 0.0001117 ecr4_3:0.00444 ecr4_3:533.333
5 263.77 290.03 0.58 0.30333 0.44167 0.0002235 0 0.0002235 0.0002235 ecr5_4:0.00742 ecr5_4:533.333
6 320.37 263.67 0.56 0.28333 0.42167 0.0002429 0 0.0002429 0.0002429 ecr6_5:0.00575 ecr6_5:533.333
7 342.18 287.59 0.27333 0.46 0.36667 0.0034641 0 0.0034641 0.0034641 ecr7_6:0.00762;ecr7_8:0.01442 ecr7_6:533.333;ecr7_8:533.333
8 246.37 247.17 0.49 0.30667 0.39833 0.0001891 0 0.0001891 0.0001891 ecr8_9:0.0054 ecr8_9:533.333
9 397.17 69.63 0.59333 0.31 0.45167 0.0006543 0 0.0006543 0.0006543
10 272.59 192.26 0.0134 0.0044 0.0089 0.0002901 0 0.0002901 0.0002901 ecr10_11:0.00676;ecr10_9:0.01087 ecr10_11:533.333;ecr10_9:533.333
11 250.45 93.71 0.45667 0.41333 0.43500 0.0001855 0 0.0001855 0.0001855
Table 1: Calibration data from the IBM quantum device ibm_sherbrooke, which was used to construct the realistic noise model. The ECR gate, used as a basis gate in IBM quantum processors, is a maximally entangling gate and is equivalent to a CNOT gate up to single-qubit pre-rotations. The numbers indicate the qubits to which the gate is applied and the order in which it is applied. Note that blank spaces for the ECR gate in some rows indicate that the gate is only available in the specific order given by the table.